Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, December 08, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 08, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/08/98 (Tuesday)

Segment 1 report for 12/8/98: At Emidio Pump Station, calibration of the pumps continued as well as installation of the control panel in the office. Backfill was in progress at the Hwy. 166 bore (EMP 1.18). The exit bore pit was excavated on the north side of the California Aqueduct. A pullbox was being installed at the Stream Crossing #4 bore (EMP 9.03). The bore operation continued at Stream Crossing #9. On Grapevine Rd., welding and x-ray took place from EMP 15.50 to 15.46. Shading took place from EMP 16.32 to 16.20. On Tejon Bench a segment of pipe was placed and tied-in to 18.07. On Lebec Rd., placement of conduit and backfill occurred from EMP's 21.46 to 21.49 and excavation occurred to EMP 21.55. Negative buoyancy pipe to be pushed through the Cuddy Creek Tributary bore (EMP 25.47) was welded, x-rayed and coated.

Segment 4B (Lynwood to El Segundo): General clean up occurred at 110th St and the Western bore site. Further, crews welded pipe in the Lynwood yard. At Lynwood station crews began construction of the block wall enclosing the site. In addition, crews trenched across little Alameda Blvd ,set duct for station power then slurried the ditch. Fiber optic crews were pulling cable from Carson St to Wilmington station. At the Tosco station, crews are working to complete the bore tie-in to the refinery.

Segment 2B (Castaic to Sylmar): Report for 12-8-98. General clean up of the right of way and removal of steel plates occurred today. In addition, x-ray and coating crews continued work south of Weldon Cyn while slurry and fiber optic duct were set nearby.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 8 December 1998: 3A: Welding on the 6" fiber optic casing pipe continued at Tuxford Street today. The main focus of activity on this segment was the completion of the tie in to Pacoima Wash Bridge crossing. Welding finished this task, with x-ray and coating crews working close behind them. Some backfill and compaction were also done in this area. Initial work on the electrical tie in to the block valve near the dog pound was performed today. 3C: Crews worked on the tie in to the Hubbard bore today, including welding, x-ray, and coating of pipe strung above ground. Other work consisted of the placement of the top of the aqueduct crossing as well as work on the high speed channel cover. Pavers worked on the area just north of the aqueduct crossing.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page